# Vacuum Rabi Splitting and Quantum Fisher Information of a Non-Hermitian Qubit in a Single-Mode Cavity

### v1

@@ III Dynamics and vacuum Rabi splitting
Figure 8: (colored online) (Upper panel) The QFI as a function of coupling strength $g$: (a) in the ground state, with $\epsilon=0.2$ (dashed line) and $\epsilon=0.4$ (solid line). The black line corresponds to the Hermitian QRM, the red line to the PTQRM, and the blue line to the symmetric QRM; (b) In the 4th excited state (black lines) and the ground state (red lines) for the Hermitian QRM (solid line) and the PTQRM (dashed line) at $\epsilon=0.2$ and $\Delta=0.5$. The QFI is scaled by $\epsilon/\epsilon_{F_{max}}$. (Lower panel) The QFI as a function of bias $\epsilon$: (c) in the ground state for $g=0.1$ (green line), $0.5$ (blue line), and $1$ (red line) in the PTQRM, and for NH-TLS (black line). Dashed lines represent the corresponding analytical results for the PTQRM (Eq.  (30)) and NH-TLS (Eq.  (29)); (d) The QFI from (c) scaled by $\epsilon/\epsilon_{EP}$, where $\epsilon_{EP}$ is the corresponding exact value for the EP.
@@
Finally, the QFI can serve as a signature for the EPs. This indicates that the quantum critically enhanced effect can locate the EPs of the PTQRM, offering an alternative to the G-function technique discussed in Sec. II(a). As illustrated in Fig. 8 (b), the QFI at $\epsilon=0.2$ for both the ground state and the excited state display sharp peaks exactly at the corresponding EPs, which align perfectly with the EPs in Fig. 2 (b).
@@ V Conclusion
We use the Bogoliubov operator approach to derive the G-function for the PTQRM, whose zeros exactly determine the eigenvalues in both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-broken regimes. We also derive the real form of the G-function, which accounts for the real eigenvalues in the $\mathcal{PT}$-unbroken phase. Using this real G-function and its first-order derivative, we can analytically detect the EPs, at which both the G-function and its first-order derivative with respect to the eigenvalues are simultaneously zero. The previous AA method can be formulated within the current BOA scheme by considering the transitions between the same manifolds. Furthermore, we improve the AA method by additionally incorporating the transitions among the nearest-neighboring manifolds. Compared to the exact solutions, the present CAA improve the eigenvalues, more importantly, the eigenstates over the AA in both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-broken regimes.
@@
We then study the VRS and atomic dynamics using the CAA in the PTQRM in various coupling regime and different imaginary biases. The emission spectrum shows two main peaks with unequal height in the ulrastrong coupling regime. The VRS becomes of multi-Rabi frequencies with increased coupling strength and higher biases. The width of the peaks also increase with the coupling strength and the imaginary biases, reflecting the nature of the open quantum systems, such as the atom-cavity coupling systems dissipating in the environment. This is due to the non-Hermitian description inheriting the nature of the Markov dynamics of the dissipative systems. Interestingly, the CAA results agree well with exact solution in all cases, indicating the CAA, where the transitions within the same manifold and between the nearest-neighboring manifolds are considered, captures the main features of the atomic dynamics and emission spectrum in this non-Hermitian light-matter interaction systems. In contrast, the AA completely fail to describe the dynamics of the this system, indicating the transition between the same manifold is insufficient to describe the dynamics. It is suggested that the present prediction for the VRS could be observed in the experiments that this model might be realized.
@@
Finally, we also calculate the QFI of the PTQRM, which indicates that quantum criticality-enhanced effects also exist in this non-Hermitian system. Interestingly, near the EPs, the QFI in the PTQRM is much higher than that in its Hermitian counterpart. More interestingly, compared to the NH-TLS, the QFI is significantly enhanced in the PTQRM in the deep-strong coupling regime. The latter is just a NH-TLS coupled with a cavity, indicating that the atom-cavity coupling can enhance the sensitivity, providing a new way to increase the estimation precision.
@@
In summary, we have proposed analytical methods to solve a class of many-body non-Hermitian quantum systems, bridging the gap between Hermitian and non-Hermitian systems, as well as between simple and complex non-Hermitian systems. We expect that the VRS could be also observed in this non-Hermitian atom-cavity coupling systems. We show that a simple non-Hermitian system, further coupled with a new degree of freedom, provides a new avenue to increase quantum sensitivity, serving as a cornerstone for future exploration.
@@ Acknowledgements.
This work is supported by the National Key R&\&&D Program of China (Grant No. 2024YFA1408900) and the National Science Foundation of China (Grants No. 12305032).
Figure 9: (colored online) The real part of the eigenvalues given by CAA and AA as a function of coupling strength $g$ at resonance ($\Delta=1$), with $\epsilon=0.1$. The left panel displays the results of $m=2$, while the right panel shows results for $m=0$. The black lines represent AA results, while the red lines denote CAA results.
@@
In the CAA, the $m$-th pair of eigenvalues  can be obtained by select the following determinant in a $6$-by-$6$ block form
@@
Note that for $m=0$, the corresponding determinant takes a $44-by-$ 4 block form, because the $m=-1$ manifold does not exist. The determinant yields six roots for $m>0$ and four roots for $m=0$, but only two are required for any $m$. As we mentioned in the main text, the CAA eigenvalues should match the AA eigenvalues. As a detailed example, Figure  9 shows six (four) groups of the real roots for $m=2(m=0)$ in the left (right) panel. Two solutions for real eigenvalues in the AA are also collected. One can immediately see that the third and the fourth roots (the first and the second roots) of the determinant fit the corresponding real eigenvalues levels of the AA for $m>0(m=0)$, while the other solutions are far from the AA results. This selection rule applies to arbitrary $m$. The corresponding imaginary parts of these two roots can also be calculated simultaneously.

### v2

@@ III Dynamics and vacuum Rabi splitting
Figure 8: Top row: (a) The difference in QFIs between the PTQRM and the Hermitian QRM, and (b) the difference between the PTQRM and the NHTLS, as a function of coupling strength $g$ and bias $\epsilon$ in the ground state. Middle row: (c) The QFI as a function of coupling strength $g$ in the ground state, with $\epsilon=0.2$ (dashed line) and $\epsilon=0.4$ (solid line). The black line represents the Hermitian QRM, the red line represents the PTQRM, and the blue line represents the symmetric QRM. The QFI is scaled by $g/g_{EP}$ or $g/g_{s}$, where $g_{EP}$ is the corresponding exact value at the EP and the $g_{s}$ is the exact value for the crossover in the Hermitian QRM. (d) The QFI as a function of bias $\epsilon$ in the ground state for $g$ = 0.1 (green line), 0.5 (blue line), and 1 (red line) in the PTQRM, and for NHTLS (black line). Dashed lines represent the corresponding analytical results for the PTQRM [Eq.  (31)] and NHTLS [Eq.  (30)]. The QFI is scaled by $\epsilon/\epsilon_{F_{max}}$, where $\epsilon_{EP}$ is the exact value at the EP. Bottom row: Time $T$ required to prepare the ground state of the systems: (e) for different biases $\epsilon$ and (f) for different coupling strengths $g$. The legends of (e) and (f) are the same as those for (c) and (d), respectively. The inset of (f) gives the QFI over the preparation time as a function of normalized biases for different coupling strengths .
Although a peak appears in the QFI near the EPs, a higher QFI does not imply directly better metrological performances. Here, we also evaluate the time required to reach the ground state toward the EPs of the PTQRM. Critical slowing down also occurs in the non-Hermitian systems, where the time required to prepare the ground state becomes extremely slow as it approaches the EPs, similar to open and closed quantum Rabi models [64, 63, 62]. We consider an adiabatic sweep where the ground state reaches the equilibrium state, representing the static approach in critical quantum metrology. We compare the time $T$ required to prepare the ground state among the PTQRM, the Hermitian QRM (HQRM), and the NHTLS. We calculate the preparation time using the bound $T\left(\lambda_{c}\right)\geq\int_{0}^{\lambda_{c}}\frac{1}{\tilde{\Delta}(\lambda)}d\lambda$ to calculate the preparation time, based on the evolution time $T=\int_{0}^{\lambda}\frac{1}{v(\lambda)}d\lambda$, where $v(\lambda)=d\lambda/dt$ is a speed at which the parameter $\lambda$ evolves during the adaptative process[64]. The condition $d\lambda/dt\ll\tilde{\Delta}(\lambda)$ holds, where the $\tilde{\Delta}(\lambda)$ is the energy gap between the ground state and first excited state[64]. In adiabatic evolution, the time required to reach the ground state is typically inversely related to the energy gap $\tilde{\Delta}(\lambda)$. As the gap closes at the EPs, the dynamical timescale $\tilde{\Delta}(\lambda)^{-1}$ diverges, causing the time required to prepare the ground state to become extremely slow. Figure 8(e) shows that the preparation time in the PTQRM is generally longer than the H-QRM. As the bias $\epsilon$ increases, which corresponds to a higher QFI in the PTQRM, the preparation time also increase. Figure 8(f) shows that the preparation time in the PTQRM is slightly longer than in the NHTLS, but a much higher QFI can be achieved by controlling the coupling strength $g$ in the PTQRM. As the non-Hermitian system approaches the EPs, the energy gap becomes very small but nonzero, and the QFI increases significantly. This implies that, despite the longer preparation time, the sensitivity of the system is still enhanced. The quantity $F/T$ is shown in the inset of Fig. 8(f). Near the EPs, the QFI increases more rapidly than the preparation time, offering an advantage in efficiently achieving higher QFI and enhanced sensitivity before reaching the EPs.
@@
Finally, the QFI can serve as a signature for the EPs. This indicates that the quantum-criticality-enhanced effect can locate the EPs of the PTQRM, offering an alternative to the $G$ function technique discussed in Sec. II A. Furthermore,the QFI for both the ground state (cf. Fig. 8 ) and the excited states (not shown here) display sharp peaks exactly at the corresponding EPs.
@@ V Conclusion
We use the Bogoliubov operator approach to derive the $G$ function for the PTQRM, whose zeros exactly determine the eigenvalues in both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-broken regimes. We also derive the real form of the $G$ function, which accounts for the real eigenvalues in the $\mathcal{PT}$-unbroken phase. Using this real $G$ function and its first-order derivative, we can analytically detect the EPs, at which both the $G$ function and its first-order derivative with respect to the eigenvalues are simultaneously zero. The previous AA method can be formulated within the current BOA scheme by considering the transitions between the same manifolds. Furthermore, we improve the AA method by additionally incorporating the transitions among the nearest-neighboring manifolds. Compared to the exact solutions, the present CAA improve the eigenvalues and, more importantly, the eigenstates over the AA in both the $\mathcal{PT}$-symmetric and $\mathcal{PT}$-broken regimes.
@@
We then study the VRS and atomic dynamics using the CAA in the PTQRM in various coupling regimes and different imaginary biases. The emission spectrum shows two main peaks with unequal height in the ulrastrong-coupling regime. The VRS becomes of multiple Rabi frequencies with increased coupling strength and higher biases. The width of the peaks also increase with the coupling strength and the imaginary biases, reflecting the nature of the open quantum systems, such as the atom-cavity coupling systems dissipating in the environment. This is due to the non-Hermitian description inheriting the nature of the Markov dynamics of the dissipative systems. Interestingly, the CAA results agree well with the exact solution in all cases, indicating the CAA, where the transitions within the same manifold and between the nearest-neighboring manifolds are considered, captures the main features of the atomic dynamics and emission spectrum in this non-Hermitian light-matter interaction system. In contrast, the AA completely fail to describe the dynamics of the system, indicating the transition between the same manifold is insufficient to describe the dynamics. The predicted VRS can be observed experimentally by realizing this model.
@@
We finally calculate the QFI of the PTQRM, which indicates that quantum criticality-enhanced effects also exist in this non-Hermitian system. Interestingly, near the EPs, the QFI in the PTQRM is much higher than that in its Hermitian counterpart. More interestingly, compared to the NHTLS, the QFI is significantly enhanced in the PTQRM in the deep-strong coupling regime. The latter is just a NHTLS coupled with a cavity, indicating that the atom-cavity coupling can enhance the sensitivity, providing a way to increase the estimation precision.
@@
In summary, we have proposed analytical methods to solve a class of many-body non-Hermitian quantum systems, bridging the gap between Hermitian and non-Hermitian systems, as well as between simple and complex non-Hermitian systems. We expect that the VRS could also be observed in these non-Hermitian atom-cavity coupling systems. We show that a simple non-Hermitian system, further coupled with a new degree of freedom, provides a avenue to enhance quantum sensitivity, serving as a cornerstone for future exploration. Future research should investigate quantum sensing using the dynamical approach, as it may provide an alternative pathway for protocol optimization, circumventing the strict dependence on precise state preparation or adiabatic conditions [65].
@@ Acknowledgements.
This work is supported by the National Key R&\&D Program of China (Grant No. 2024YFA1408900) and the National Science Foundation of China (Grant No. 12305032).
Figure 9: The real part of the eigenvalues given by CAA and AA as a function of coupling strength $g$ at resonance ($\Delta=1$), with $\epsilon=0.1$: (left) results for $m=2$ and (right) results for $m=0$. The black lines represent AA results, while the red lines denote CAA results.
@@
In the CAA, the $m$-th pair of eigenvalues can be obtained by select the following determinant in a $6\times 6$ block form
@@
Note that for $m=0$, the corresponding determinant takes a $4\times 4$ block form, because the m=1m=-1 manifold does not exist. The determinant yields six roots for $m>0$ and four roots for $m=0$, but only two are required for any $m$. As we mentioned in the main text, the CAA eigenvalues should match the AA eigenvalues. As a detailed example, Fig.  9 shows six (four) groups of the real roots for $m=2(m=0)$ in the left-hand (right-hand) panel. Two solutions for real eigenvalues in the AA are also collected. One can immediately see that the third and the fourth roots (the first and the second roots) of the determinant fit the corresponding real eigenvalues levels of the AA for $m>0(m=0)$, while the other solutions are far from the AA results. This selection rule applies to arbitrary $m$. The corresponding imaginary parts of these two roots can also be calculated simultaneously.